OpenStudy (anonymous):

8.) Solve the following system of equations y + 4x = 8 and 5x + 2y = 13 (A) 1, 4********* (B) 3/2, 1/2 (C) -1/2, 4

OpenStudy (anonymous):

Substitute the first equation for Y in the second equation...

OpenStudy (anonymous):

5x+2(-4x+8)=13

OpenStudy (phi):

yes. Though you can use substitution, elimination (especially if you do more than 2 variables) is preferable. in the order Ax+By= constant we have 4x +1 y= 8 , and 5x + 2y = 13 dropping x,y 4 1 8 5 2 13 multiply the top row by -5/4 and add to the 2nd row 4 1 8 0 3/4 3 backsolve 3/4 y = 3 y= 4 use the first row, with y=4 4x+4= 8 x=1

OpenStudy (anonymous):

Thank you. It is the right answer

OpenStudy (anonymous):

Thank you. It is the right answer

OpenStudy (anonymous):

Thank you. It is the right answer